OpenStudy (anonymous):

A= h(m+t) ------- solve for M 2 ? please show me some work so i understand, thanks

OpenStudy (anonymous):

is it 2A-ht = m ---- h ?

OpenStudy (anonymous):

it is ((2A)/h)-t=M

OpenStudy (anonymous):

2A=h(m+t) 2A/h=m+t (2A/h)-t=M

OpenStudy (anonymous):

Did u get that?

OpenStudy (anonymous):

i M= 2A-ht _______ h

OpenStudy (anonymous):

i mean I got M= 2A-ht _______ h

OpenStudy (anonymous):

ya but separate the answer like this

OpenStudy (anonymous):

M = (2A/h)-(ht/h)

OpenStudy (anonymous):

i just sepaprated the equation. now do u see that u can cross out the h for ht/h

OpenStudy (anonymous):

so it's 2a-t=m ?

OpenStudy (anonymous):

ummm no

OpenStudy (anonymous):

2A/h - ht/h

OpenStudy (anonymous):